Consider the following reaction: Bra(g) + Cla(g) 2 BrCl(g) After 0.60 mol of Brz and 0.60 mol of Ch were placed in a 1.00 L flask, the reaction was allowed to reach equilibrium. After reaching equilibrium, the flask is found to contain 0.28 mol of BrCl. What is the value of K for this reaction?
1
Expert's answer
2020-05-08T14:12:05-0400

Br2(g) + Cl2(g) = 2BrCl(g)


K = xBrCl2/(xBr2*xCl2), where x - molar fraction of component in the mixture.


Find the number of moles of components at equilibrium:


nequilibrium(Br2) = ninitial(Br2) - nreact(Br2) = ninitial(Br2) - nreact(BrCl)/2 = 0.60 - 0.28/2 = 0.46 mol

nequilibrium(Cl2) = ninitial(Cl2) - nreact(Cl2) = ninitial(Cl2) - nreact(BrCl)/2 = 0.60 - 0.28/2 = 0.46 mol


Calculate the molar fractions for each component (x = ncomponent/nmixture):


xBrCl = n(BrCl)/nmixture = 0.28/1.2 = 0.233

xBr2 = nequilibrium(Br2)/nmixture = 0.46/1.2 = 0.383

xCl2 = nequilibrium(Cl2)/nmixture = 0.46/1.2 = 0.383


Сalculate K for this reaction:

K = xBrCl2/(xBr2*xCl2) = 0.2332/(0.383*0.383) = 0.37


Answer: K for reaction Br2(g) + Cl2(g) = 2BrCl(g) is 0.37.